Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
Technical integration and coordination SASE1 Status and SASE3 Readiness
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 591     Entry time: 26 Feb 2025, 09:24
Author: Andreas Galler 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 26.02.2025 

General Announcement:
-     

CTRL:
- nothing major

DA:
- nothing major

EEE:
- PLC large detector mover worked yesterday but not on the weekend (MID)
- CRL1 Lens2Y SA1 encoder not working
   

VAC:
- only minor thing

XPD:
- ZZ yesterday
   - pre vacuum for PES in XTD10, solved quickly and things back to normal. --> a valve was not completely closed putting load on the scroll pump thus in

DET:
- Nothing major

ITDM: 
- nothing to report

XRO:
- access to XTD2
  - exchanging proximity sensor for lens1y in CRL1, XTD2
  - checked SRA on issues

LAS:
- not to report

UNSYS:
- nothing to report

SPB:
- nothing to report

FXE:
- CRL5Y motor issue in hutch, solved with help of Moises, tolerance had to be tweaked.
 

MID:
- all fine and thanks to EEE for helping with the large detector mover.

HED:
- all good

SCS:
- nothing to report

SQS:
- not present

SXP:
- nohting, all set

ELOG V3.1.4-7c3fd00